the ionisation potential  and electron affinity of an element is 275 &86 kcal per mole then EN of the element inpaulings scale will be 
 

Grade:12th pass

2 Answers

Arun
4 years ago
Xm=I.P+E.A/125 (In kcal/mol this formula is used bcz energies are given in kcal/mol) 
So ans = 275+86/125 = 2.888
Vikas TU
4 years ago
Dear student 
Xmuliken = (IE+IA)/2 
IE = Ioniation enthalpy  = 275 kcal/mole
EA = electron affinity = 86 kcal/mole
Xpauling = 0.33 (Xmilliken - 0.615)
X mulliken = 180.5 kcal /mol
Xpauling  0.336(180.5 - 0.615) = 60.44 Kcal/mol
Therefore, the electronegativity of element in Pauling scale will be, 60.44 kcal/mole